Weekly LIPAc highlights (10/May-14/May)


The injector conditioning using 6mm Plasma Electrode is currently on-hold waiting for the procurement and repair of the cooling plate of the Power Supply of the Solenoid 1. The hardware delivery is planned by last week of June.

Radio Frequency (RF) System and RFQ cavity

In the last week the accelerator vault was open in the day time for the on-going activities on the HEBT (High Energy Beam Transport) and BD (Beam Dump).

RF injection in the RFQ cavity was resumed daily after 5PM when the vault could be closed. During the night the RF-RFQ injection was monitored by NAT operator and injection was mostly stable at around 5% duty cycle, without major disruptions. RF team checks during the injection focused on the stability, noise behaviour and circulator controllability.

The repair of the flowmeter of the cooling line of one of two Solid State Power Amplifiers (SSPA) has been completed.

Beam Dump

The alternative chemical for pH control (AM-490) was delivered and the first circulation tests with oxygen control (10 ppb is not reached yet) and pH control (using AM-490) were done with remote participation from CIEMAT.

The connection of the cooling loop to BD cartridge was completed in the vault (a temporary by-pass in front of the cartridge was installed so far) and filling of the entire loop including the cartridge is currently on-going. Circulation tests including the cartridge are planned starting form next week.

Control System

Connectivity checks for MPS/PPS (Machine Protection System/Personal Protection system) with HEBT/BD LCS (Local Control System) are on-going. Efforts focused on the Lead shutter.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.